OpenStudy (et365):

Integral from 0,4 (5-2x)dx, guys whats the simplest way to work this??

Vocaloid (vocaloid):

well, do you know how to calculate the antiderivative (integral) of 5 - 2x?

OpenStudy (et365):

That would be x^2+c correct?

OpenStudy (solomonzelman):

No

OpenStudy (et365):

or rather 5-x^2|0,2?

OpenStudy (solomonzelman):

I will give you an example that is similar to this: \(\large\color{black}{ \displaystyle \int_{ 2 }^{8}\left(7-10x\right){~}dx }\) *[1]* An integral of a constant *b* (with respect to x) is equal to *bx*. (For any non-zero number *b*.) *[2]* An integral of any \(x^n\) (with respect to x) is \(x^{n+1}/(n+1)\), and integral of b•\(x^n\) (when you have a coefficient b in front) is \({\rm b}x^{n+1}/(n+1)\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ So you first integrate each term: 1) The integral of 7 is 7x (based on *[1]* ). 2) The integral of -10x (which is same as -10x\(^1\), is \({\rm -10}x^{1+1}/(1+1)=-10x^2/2=-5x^2\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ So you get: \(\large\color{black}{ \displaystyle \int_{ 2 }^{8}\left(7-10x\right){~}dx =7x-5x^2 }\) and then evaluate that from x=2 to x=8 (those are the limits of integration) In other words: \(\large\color{black}{ \displaystyle \int_{ 2 }^{8}\left(7-10x\right){~}dx =\left[7(8)-5(8)^2\right] -\left[7(2)-5(2)^2\right]=-258 }\)

OpenStudy (solomonzelman):

This is an example, read through it and ask if you have any questions.

OpenStudy (et365):

So what happens the the dx?

OpenStudy (solomonzelman):

So the \(\displaystyle \int_{a}^{b}\) and \(\rm dx\) are just indicating that you are integrating with respect to x, from x=a to x=b.

OpenStudy (solomonzelman):